# A comparative high-resolution spectroscopic analysis of in situ and accreted globular clusters

### v1

@@ 5.1 NGC 6218, NGC 6522, NGC 6626 and the Milky Way
We use as a comparison the results obtained from high resolution spectroscopy of MW field stars, focusing on the chemical elements that show significant differences in accreted and in situ GCs. To do so, given the extremely high-quality of their stellar spectra, we selected the catalogue by Nissen & Schuster (2010, 2011) as our benchmark comparison dataset, and in particular we focused on stars in the high- sequence that have been observed with UVES at VLT. To maintain homogeneity in the chemical analysis, we re-derived abundances from spectra of these stars (see details of the method in Section 3). The only exception in this homogeneous comparison is the element EuII, as (i) the number of available spectra in the literature is limited (and many of them were obtained with different instruments and S/N), and (ii) abundances are derived from different spectral lines, due to the different evolutionary stages of the observed stars. We choose to compare our results with EuII abundances from the high- sequence of Fishlock et al. (2017).
@@
In Figure 7 we show that the abundances of NGC 6218, NGC 6522, and NGC 6626 match that of field stars (gray filled points) in all chemical spaces of interest (i.e., Ca, TiI, Zn, YII and EuII) except for Mg, where the three in situ target GCs show enhancement in [Mg/Fe] compared to MW high- sequence stars, being outside the of their distribution in the metallicity range of .
5.2 NGC 362, NGC 1261 and Gaia-Sausage-Enceladus
@@
The reference catalogue used for this comparison is the one by Ceccarelli et al. (2024), which has been produced with the same methods described here. Additionally, we re-derived the abundances for the stars in the low- sequence of Nissen & Schuster (2010). Finally, we compared our abundances with the results from the low- stars in Fishlock et al. (2017), Aguado et al. (2021), Carrillo et al. (2022), Giribaldi & Smiljanic (2023) and François et al. (2024) for EuII. As displayed in Figure 8, there is excellent agreement between the chemical composition of NGC 362 and NGC 1261 and the patterns shown by GSE stars (gray filled points) in the -elements Ca and TiI, while, as for in situ GCs, the two GCs lay in the upper boundaries (at the 1.3 and 1.6 levels for NGC 362 and NGC 1261, respectively) of the distribution of GSE field stars in [Mg/Fe] in the metallicity range of interest. The value of [Zn/Fe] derived in NGC 1261 is consistent with those observed in the most Zn-poor stars in GSE, while NGC 362 seems to be slightly depleted () compared to them. Finally, the slightly subsolar values of [YII/Fe] are consistent with those observed in literature (Aguado et al. 2021, Ceccarelli et al. 2024, François et al. 2024), as well as the high [EuII/Fe] ratios (see also, Aguado et al. 2021, Matsuno et al. 2021, François et al. 2024, Ou et al. 2024).
@@ 5.3 Ruprecht 106 and the Helmi Streams
In this Section we investigate the compatibility of the chemical composition of Rup106 and that of stars formed in the galaxy that originated the Helmi Streams (Helmi et al. 1999), as a tentative dynamical connection between the two has been pointed out by several works (Massari et al. 2019, Forbes 2020, Callingham et al. 2022). In Figure 9, we present the mean abundances of Rup106 derived in this work (blue filled squares) compared to the high-resolution abundances of 11 stars linked to the Helmi Streams (grey filled points, Matsuno et al. 2022a, M22 hereafter). Note that we cannot compare Eu abundances as they were not published in M22, and we thus show Ni as an additional comparison, representative of the iron-peak group. To ensure homogeneity in the comparison, we re-analyze the spectra of the Helmi Streams stars following the procedure described in Section 3. The results are shown in Figure 9 (black filled points). The average offset between our and M22 abundances is consistent with zero for all the elements we show in Figure 9, except for TiI, Ni and YII, for which we find mild offsets of (), of (), and of (), respectively. Regardless of the adopted data set (even though the most homogeneous comparison is the one with our re-analysed abundances), when comparing the mean abundance ratios of Rup106 to the trend observed in the Helmi Streams, a noticeable difference emerges, particularly in the -elements. Our findings consistently indicate values approximately lower in comparison to the Helmi Streams, hinting at a lower star formation efficiency in the environment where Rup106 formed. Among the other elements studied, the most significant differences arise in Ni and YII with Rup106 being depleted by and by respectively. Therefore, in order to account for the observed differences shown in Figure 9, the progenitor of the Helmi Streams should have been massive enough to be chemically inhomogeneous. Otherwise, this chemical incongruity is a strong evidence supporting the interpretation that Rup106 was born in a different environment compared to the progenitor of the Helmi Streams.
6 On the possible origin of Ruprecht 106
In light of the chemical difference between Rup106 and the Helmi Streams stars, as well as because of its peculiar chemical composition compared to many other GCs (e.g., Brown et al. 1997, Carretta et al. 2009, 2013, Villanova et al. 2013, Monaco et al. 2018, Puls et al. 2018, Crestani et al. 2019, Masseron et al. 2019, Horta et al. 2020, Mucciarelli et al. 2021b), the birth environment of Rup106 likely underwent a unique chemical evolution. In the following we discuss some possible interpretations.
@@ 6.1 Comparison with NGC 121
Given its extremely peculiar chemical composition, Rup106 has long been regarded as originating in an extragalactic environment. Initially, several work proposed it to be accreted from the Magellanic Clouds (Lin & Richer 1992, Fusi Pecci et al. 1995), and particularly from the Small Magellanic Cloud (SMC). Even though recent post-Gaia kinematics studies of the orbit in the MW of Rup106 (see Section 1) have excluded the possibility of it originating in the SMC, it is interesting to note that Rup106 is coeval with NGC 121, the oldest SMC globular cluster ( Gyr, Glatt et al. 2008, Forbes & Bridges 2010, Dotter et al. 2011) and shows very similar metallicity (Dalessandro et al. 2016, Mucciarelli et al. 2023). Therefore, this presents an excellent opportunity to explore the possibility of having an SMC-like galaxy as progenitor of Rup106. We will directly compare our results with Mucciarelli et al. (2023), given the congruent analytical approaches. As shown in Figure 10, NGC 121 (yellow filled squares) has solar-scaled , hinting at an enrichment due to SN Ia already in place in the SMC Gyr ago. These values are, on average, - -enriched relative to what we find for Rup106, suggesting that Rup106 likely formed in an environment with an even less efficient star formation compared to the SMC. Regarding the iron-peak elements, we find comparable values in the abundances ratios of Cu. The very low values of [Cu/Fe] found in both Rup106 and NGC 121 are consistent with those of the SMC field stars (Mucciarelli et al. 2023) and 0.5 dex underabundant compared to MW GCs and stars, favouring a chemical enrichment scenario in which the contribution by massive stars is extremely reduced, as also discussed in Section 4.2. Additionally, Rup106 shows lower abundances () of other elements produced in massive stars (i.e., Sc, V, and Co) relative to NGC 121 and also in Ni. Interestingly, the only elements in which enhancement is detected in Rup106 are Cr and Mn. These elements are predominantly synthesized by SN Ia, with yields depending on the metallicity of the progenitor SN Ia. Unfortunately, Mucciarelli et al. (2023) do not provide an estimate of [Zn/Fe] for NGC 121. When assessing -process elements, we note similar Ba abundances between the two clusters, yet discern distinct values in Y and La, with Rup106 being, once again, depleted in these elements. Given that NGC 121 demonstrates pronounced deficiencies in these elements relative to the only other SMC GC of similar age, this discrepancy highlights the unique conditions under which Rup106 formed. Also, the efficiency of production of -process element is extremely low, with Rup106 being underabundant in Eu relative to NGC 121, which conversely aligns with observations in GSE GCs and SMC stars. This comparison further underlines that Rup106 originated in an environment that was significantly different to the SMC of Gyr ago, probably characterized by an exceptionally low star formation efficiency and likely featured a truncated initial mass function (IMF) with minimal contribution from massive stars.
@@ 6.2 Comparison with surviving dwarf galaxies
A similar chemical composition to that of Rup106 has been observed in stars belonging to dwarf galaxies in the Local Group (Venn et al. 2004, Tolstoy et al. 2009). Indeed, subsolar values at are observed in surviving dwarf galaxies like Carina, Fornax, Sculptor, Sextans, and Ursa Minor (Venn et al. 2012, Hendricks et al. 2014, Hill et al. 2019, Theler et al. 2020, Fernandes et al. 2023, Sestito et al. 2023). The low values of Zn ([Zn/Fe] ) compared to MW stars, as derived for Rup106, are also typical these systems (see Fig. 16 of Skúladóttir et al. 2017, and reference therein). Finally, most of the dwarf galaxies show enhancement in [EuII/Fe] compared to MW stars due to either the very low star formation efficiency and the delayed contribution to the production of Eu by NSMs (Letarte et al. 2010, Venn et al. 2012, Lemasle et al. 2014). However, even the extremely low [EuII/Fe] of Rup106 can be reproduced in some of these systems, like, e.g., Sculptor (Hill et al. 2019). Indeed, these authors find several stars in Sculptor with subsolar values () of [EuII/BaII] at , suggesting that the onset of -process enrichment was already in place at this metallicity. It is interesting to note that the chemical composition of Rup106 recalls that of iron-rich metal-poor stars (IRMP), despite they being unlikely to form in GCs, thus indicating a birth environment enriched by the (double) detonation of sub-Chandrasekhar-mass CO white dwarfs (Reggiani et al. 2023). This further support the idea that Rup106 formed in a dwarf galaxy, as these authors demonstrated that IRMP stars are easier to find in surviving dwarf galaxies and the Magellanic Clouds, rather than in the halo of the MW.
It is important to note that the majority of the existing systems that reproduce the peculiar chemistry of Rup106 have a stellar mass ( , McConnachie 2012) that is not compatible with the one that has been estimated for the progenitor of the Helmi Streams, with the latter being roughly 100 times more massive (Koppelman et al. 2019). Therefore, if the dynamical association of Rup106 is accurate, the possibility that it did not form directly in the progenitor of the Helmi Streams cannot be ruled out. We speculate that Rup106 could have formed in a dwarf galaxy with a mass comparable to that of Sculptor, that, based on Rup106 dynamical properties, could have been a satellite of the progenitor of the Helmi Streams or the Sagittarius dwarf (Davies et al. 2024), on a similar note to the case of the GC NGC 2005 in the Large Magellanic Cloud (Mucciarelli et al. 2021b).
@@ 7 Conclusion
The primary aim of this work is to discern potential distinctions in elemental abundances among four GCs of very similar metallicity that are expected to have formed in different environments, and to understand which elements are most sensitive to their different origin.
@@
The outcome of this homogeneous chemical analysis reveals striking similarities in the chemical abundances of NGC 362 and NGC 1261, two GCs that were accreted during the merging event with the Gaia-Sausage-Enceladus dwarf galaxy. When comparing their abundances with those from the in situ GCs NGC 6218, NGC 6522 and NGC 6626, a coherent and distinct overall pattern emerges in the -elements, especially in Mg, Si, and Ca, where we identify an depletion of in the GSE GCs with respect to in situ GCs. This finding agrees with observations of Halo field stars, where distinctions between in-situ and accreted stars, especially in the -elements, are discernible around this values of [Fe/H], whereas chemical patterns become indistinguishable at lower metallicity (Horta et al. 2023, Ceccarelli et al. 2024). This is consistent with the fact that it took 2 Gyr less to NGC 6218, NGC 6522 and NGC 6626 to reach the same metallicty as the GSE GCs (Dotter et al. 2010, 2011, VandenBerg et al. 2013, Villanova et al. 2017, Kerber et al. 2018), indicating that the contribution of SN Ia to the chemical enrichment of the gas was not in place at the time when this system formed. Moreover, statistically significant variations are observed in specific elements such as Zn and Eu, whose production sites are directly associated with massive stars (HNe for Zn and rCCSNe+NSMs for Eu). It is noteworthy to emphasize that these events are inherently rare, thus they might exert a profound influence on the abundances within a given environment. These results reinforce the idea that chemical tagging is an effective tool to investigate the origin of GCs, even when comparing chemical composition of GCs formed in different progenitor, but at the same time warn that not every element is equally sensitive in this sense.
@@
In this coherent picture, Rup106 stands out as a very distinct character. Indeed, Rup106 follows the AMR of accreted GCs (Forbes & Bridges 2010, Dotter et al. 2010, 2011, VandenBerg et al. 2013), whereas MW in situ clusters of similar age clearly experienced a different chemical enrichment path, exhibiting significantly higher metallicity (0.5 - 0.8 dex more metal rich, Massari et al. 2023). Despite the congruence between Rup106 and the GSE GCs in the age-metallicity space, differences emerge when examining all their chemical elements. Particularly, Rup106 exhibits subsolar values, that are typical of an environment already enriched by SNe Ia, in contrast to GSE clusters (), where SNe Ia have only just started contributing to the pollution of the gas. This comparison underlines distinct chemical evolution histories between GSE and the environment in which Rup106 formed, with the latter experiencing slower and less efficient star formation and, likely, an IMF truncated of most of the massive stars, indicating a lower mass for the host of Rup106 at 10.5 Gyr ago compared to GSE.
@@
In light of these results, this work demonstrates that the exceptionally high precision offered by a strictly homogeneous chemical analysis allows chemical tagging to play a crucial role in uncovering the true origin of GCs. Particularly, this method enables to clearly discriminate not only between GCs formed within the MW and those formed in other accreted galaxies, but also to distinguish the chemical signatures of GCs originating from independent accreted progenitors.
@@
Finally, we interpret the extreme peculiarity of the chemistry of Rup106 as due to a particular evolution of its progenitor, possibly the galaxy originating the Helmi Streams. However, the remarkably low values in the abundance ratios, as well as some other heavier elements, compared to Helmi Streams stars lend weight to the possibility that Rup106 originated from an even smaller galaxy, with a stellar mass comparable to that of some surviving dwarf galaxies (e.g., Sculptor).

### v2

@@ 5.1 NGC 6218, NGC 6522, NGC 6626, and the Milky Way
We use as a comparison the results obtained from high-resolution spectroscopy of MW field stars, focusing on the chemical elements that show significant differences between accreted and in situ GCs. To do so, given the extremely high quality of their stellar spectra, we selected the catalogue of Nissen & Schuster (2010, 2011) as our benchmark comparison dataset, and in particular we focus on stars in the high- sequence that have been observed with UVES at the VLT. To maintain homogeneity in the chemical analysis, we re-derived abundances from the spectra of these stars (see the detailed method in Sect. 3). The only exception in this homogeneous comparison is the element EuII, as (i) the number of available spectra in the literature is limited (and many of them were obtained with different instruments and S/N), and (ii) abundances are derived from different spectral lines due to the different evolutionary stages of the observed stars. We chose to compare our results with EuII abundances from the high- sequence of Fishlock et al. (2017).
@@
Fig. 7 shows that the abundances of NGC 6218, NGC 6522, and NGC 6626 match those of the field stars (grey filled points) in all chemical spaces of interest (i.e. Ca, TiI, Zn, YII, and EuII) except for Mg, where the three in situ target GCs show enhancement in [Mg/Fe] compared to MW high- sequence stars, which is found to be outside the of their distribution in the metallicity range of .
5.2 NGC 362, NGC 1261, and Gaia-Sausage-Enceladus
@@
The reference catalogue used for this comparison is that of Ceccarelli et al. (2024), which was produced with the same methods as those described here. Additionally, we re-derived the abundances for the stars in the low- sequence of Nissen & Schuster (2010). Finally, we compared our abundances with the results from the low- stars in Fishlock et al. (2017), Aguado et al. (2021), Carrillo et al. (2022), Giribaldi & Smiljanic (2023), and François et al. (2024) for EuII. As displayed in Fig. 8, there is excellent agreement between the chemical composition of NGC 362 and NGC 1261 and the patterns shown by GSE stars (grey filled points) in the -elements Ca and TiI, while, as for in situ GCs, NGC 362 and NGC 1261 lay in the upper boundaries (at the 1.3 and 1.6 levels for NGC 362 and NGC 1261, respectively) of the distribution of GSE field stars in [Mg/Fe] in the metallicity range of interest. The value of [Zn/Fe] derived in NGC 1261 is consistent with those observed in the most Zn-poor stars in GSE, while NGC 362 seems to be slightly depleted () compared to them. Finally, the slightly subsolar values of [YII/Fe] found for NGC 362 and NGC 1261 are consistent with those observed in the literature (Aguado et al. 2021, Ceccarelli et al. 2024, François et al. 2024), as are the high [EuII/Fe] ratios (see also, Aguado et al. 2021, Matsuno et al. 2021, François et al. 2024, Ou et al. 2024).
@@ 5.3 Ruprecht 106 and the Helmi Streams
In this section, we investigate the compatibility of the chemical composition of Rup106 and that of stars formed in the galaxy that originated the Helmi Streams (Helmi et al. 1999), as a tentative dynamical connection between the two has been pointed out by several works (Massari et al. 2019, Forbes 2020, Callingham et al. 2022). In Fig. 9 we present the mean abundances of Rup106 derived in this work (blue filled squares) compared to the high-resolution abundances of 11 stars linked to the Helmi Streams (grey filled points; Matsuno et al. 2022a, M22 hereafter). We note that we cannot compare EuII abundances as they were not published in M22, and we therefore show Ni as an additional comparison, representative of the iron-peak group. To ensure homogeneity in the comparison, we reanalysed the spectra of the Helmi Streams stars following the procedure described in Sect. 3. The results are shown in Fig. 9 (black filled points). The average offset between our abundances and those of M22 is consistent with zero for all the elements we show in Fig. 9, except for TiI, Ni, and YII, for which we find mild offsets of (), (), and (), respectively. Regardless of the adopted data set (even though the most homogeneous comparison is the one with our reanalysed abundances), when comparing the mean abundance ratios of Rup106 to the trend observed in the Helmi Streams, a noticeable difference emerges, particularly in the -elements. Our findings consistently indicate values approximately lower in comparison to the Helmi Streams, hinting at a lower star formation efficiency in the environment where Rup106 formed. Among the other elements studied, the most significant differences arise in Ni and YII with Rup106 being depleted by and by respectively. Therefore, in order to account for the observed differences shown in Fig. 9, the progenitor of the Helmi Streams would need to have been massive enough to be chemically inhomogeneous. Otherwise, this chemical incongruity is strong evidence in support of the interpretation that Rup106 was born in a different environment from the progenitor of the Helmi Streams.
6 Possible origins of Ruprecht 106
In light of the chemical difference between Rup106 and the Helmi Streams stars, and given its peculiar chemical composition compared to many other GCs (e.g. Brown et al. 1997, Carretta et al. 2009, 2013, Villanova et al. 2013, Monaco et al. 2018, Puls et al. 2018, Crestani et al. 2019, Masseron et al. 2019, Horta et al. 2020, Mucciarelli et al. 2021b), the birth environment of Rup106 likely underwent a unique chemical evolution. In the following, we discuss some possible interpretations.
@@ 6.1 Comparison with NGC 121
Given its extremely peculiar chemical makeup, Rup106 has long been regarded as originating in an extragalactic environment. Initially, several works proposed that it was likely to have been accreted from the Magellanic Clouds (Lin & Richer 1992, Fusi Pecci et al. 1995), and particularly from the Small Magellanic Cloud (SMC). Although recent post-Gaia kinematics studies of the orbit in the MW of Rup106 (see Sect. 1) have excluded the possibility that it originated in the SMC, it is interesting to note that Rup106 is coeval with NGC 121, the oldest known SMC GC ( Gyr, Glatt et al. 2008, Forbes & Bridges 2010, Dotter et al. 2011) and shows very similar metallicity (Dalessandro et al. 2016, Mucciarelli et al. 2023). Therefore, this presents an excellent opportunity to explore the possibility that an SMC-like galaxy is the progenitor of Rup106. We directly compared our results with those from Mucciarelli et al. (2023), given the congruent analytical approaches. As shown in Fig. 10, NGC 121 (yellow filled squares) has solar-scaled , hinting at an enrichment due to SN Ia already in place in the SMC Gyr ago. These values are, on average, - -enriched relative to what we find for Rup106, suggesting that Rup106 likely formed in an environment with even less efficient star formation than that of the SMC. Regarding the iron-peak elements, we find comparable values in the abundance ratios of Cu. The very low values of [Cu/Fe] found in both Rup106 and NGC 121 are consistent with those of the SMC field stars (Mucciarelli et al. 2023) and are 0.5 dex underabundant compared to MW GCs and stars, favouring a chemical-enrichment scenario in which the contribution by massive stars is extremely reduced, as also discussed in Sect. 4.2. Additionally, Rup106 shows lower abundances () of other elements produced in massive stars (i.e. Sc, V, and Co) relative to NGC 121 and also in Ni. Interestingly, the only elements in which enhancement is detected in Rup106 are Cr and Mn. These elements are predominantly synthesised by SN Ia, with yields depending on the metallicity of the progenitor SN Ia. Unfortunately, Mucciarelli et al. (2023) do not provide an estimate of [Zn/Fe] for NGC 121. When assessing -process elements, we note similar Ba abundances between the two clusters, yet discern distinct values in Y and La, with Rup106 being, once again, depleted in these elements. Given that NGC 121 demonstrates pronounced deficiencies in these elements relative to the only other SMC GC of similar age, this discrepancy highlights the unique conditions under which Rup106 formed. Also, the efficiency of the production of -process elements is extremely low, with Rup106 being underabundant in Eu relative to NGC 121, which conversely aligns with observations of GSE GCs and SMC stars. This comparison further supports the idea that Rup106 originated in an environment that was significantly different from the SMC of Gyr ago and was probably characterised by an exceptionally low star formation efficiency and likely featured a truncated initial mass function (IMF) with a minimal contribution from massive stars.
@@ 6.2 Comparison with surviving dwarf galaxies
A similar chemical composition to that of Rup106 has been observed in stars belonging to dwarf galaxies in the Local Group (Venn et al. 2004, Tolstoy et al. 2009). Indeed, subsolar values at are observed in surviving dwarf galaxies such as Carina, Fornax, Sculptor, Sextans, and Ursa Minor (Venn et al. 2012, Hendricks et al. 2014, Hill et al. 2019, Theler et al. 2020, Fernandes et al. 2023, Sestito et al. 2023). The low values of Zn ([Zn/Fe] ) derived for Rup106 compared to those found in MW stars are also typical of these systems (see Fig. 16 of Skúladóttir et al. 2017, and reference therein). Finally, most of the dwarf galaxies show enhancement in [EuII/Fe] compared to MW stars due to either the very low star formation efficiency or the delayed contribution to the production of Eu by NSMs (Letarte et al. 2010, Venn et al. 2012, Lemasle et al. 2014). However, even the extremely low [EuII/Fe] of Rup106 can be reproduced in some of these systems, such as Sculptor (Hill et al. 2019). Indeed, these latter authors find several stars in Sculptor with subsolar values () of [EuII/BaII] at , suggesting that the onset of -process enrichment was already in place at this metallicity. It is interesting to note that the chemical composition of Rup106 resembles that found in iron-rich metal-poor stars (IRMP), despite these latter being unlikely to form in GCs, thus indicating a birth environment enriched by the (double) detonation of sub-Chandrasekhar-mass CO white dwarfs (Reggiani et al. 2023). This further supports the idea that Rup106 formed in a dwarf galaxy, as these authors demonstrated that IRMP stars are easier to find in surviving dwarf galaxies and the Magellanic Clouds, rather than in the halo of the MW.
It is important to note that the majority of the existing systems that reproduce the peculiar chemistry of Rup106 have a stellar mass ( , McConnachie 2012) that is not compatible with that estimated for the progenitor of the Helmi Streams, with the latter being roughly 100 times more massive (Koppelman et al. 2019). Therefore, if the dynamical association of Rup106 is accurate, the possibility that it did not form directly in the progenitor of the Helmi Streams cannot be ruled out. We speculate that Rup106 could have formed in a dwarf galaxy with a mass comparable to that of Sculptor; the dynamical properties of Rup106 suggest that this could have been a satellite of the progenitor of the Helmi Streams or the Sagittarius dwarf (Davies et al. 2024), similarly to the case of the GC NGC 2005 in the Large Magellanic Cloud (Mucciarelli et al. 2021b).
@@ 7 Conclusion
The primary aim of this work is to investigate the potential differences in elemental abundances between six GCs of very similar metallicity that are expected to have formed in different environments, and to decipher which abundance ratios are most sensitive to their different origin.
@@
The outcome of this homogeneous chemical analysis reveals striking similarities in the chemical abundances of NGC 362 and NGC 1261, two GCs that were accreted during the merger event of the Gaia-Sausage-Enceladus dwarf galaxy. When comparing the chemical abundances of these GCs with those of the in situ GCs NGC 6218, NGC 6522, and NGC 6626, a coherent and distinct overall pattern emerges in the -elements, especially in Mg, Si, and Ca, where we identify a depletion of in the GSE GCs with respect to in situ GCs. This finding agrees with observations of field stars of the Galactic halo in this metallicity range, where distinctions between chemical patterns of in situ and accreted stars are discernible, especially in the -elements, whereas trends become indistinguishable at lower metallicity (Horta et al. 2023, Ceccarelli et al. 2024). This is consistent with the fact that it took 2 Gyr less for NGC 6218, NGC 6522, and NGC 6626 to reach the same metallicity as the GSE GCs (Dotter et al. 2010, 2011, VandenBerg et al. 2013, Villanova et al. 2017, Kerber et al. 2018), indicating that the contribution of SN Ia to the chemical enrichment of the gas was not in place at the time when these systems formed. Moreover, statistically significant variations are observed in specific elements such as Zn and Eu, whose production sites are directly associated with massive stars (HNe for Zn and rCCSNe+NSMs for Eu). It is worth noting that that these events are inherently rare, and might exert a profound influence on the abundances within a given environment. These results reinforce the idea that chemical tagging is an effective tool for investigating the origin of GCs, even when comparing the chemical compositions of GCs formed in different progenitors, but at the same time demonstrate that not every element is equally sensitive in this sense.
@@
In this coherent picture, Rup106 stands out as a highly distinct GC. Indeed, Rup106 follows the AMR of accreted GCs (Forbes & Bridges 2010, Dotter et al. 2010, 2011, VandenBerg et al. 2013), whereas MW in situ clusters of similar age clearly experienced a different chemical enrichment path, exhibiting significantly higher metallicity (0.5 - 0.8 dex more metal rich; Massari et al. 2023). Despite the congruence between Rup106 and the GSE GCs in the age–metallicity space, differences emerge when examining all their chemical elements. Particularly, Rup106 exhibits subsolar values typical of an environment already enriched by SNe Ia, in contrast to GSE clusters (), where SNe Ia have only just started contributing to the enrichment of the gas. This comparison underlines distinct chemical evolution histories between GSE and the environment in which Rup106 formed, with the latter experiencing slower and less efficient star formation and likely an IMF truncated of most of the massive stars, indicating a lower mass for the host of Rup106 at 10.5 Gyr ago compared to GSE.
@@
In light of these results, this work demonstrates that the exceptionally high precision offered by a strictly homogeneous chemical analysis allows chemical tagging to play a crucial role in uncovering the true origin of GCs. Particularly, this method enables us not only to clearly discriminate between GCs formed within the MW and those formed in other accreted galaxies, but also to distinguish the chemical signatures of GCs originating from independent accreted progenitors.
@@
Finally, we interpret the extreme peculiarity of the chemistry of Rup106 as due to the particular evolution of its progenitor, which could be the galaxy that gave rise to the Helmi Streams. However, the remarkably low values of the abundance ratios, as well as of some other heavier elements, compared to Helmi Streams stars lend weight to the possibility that Rup106 originated from an even smaller galaxy, with a stellar mass comparable to that of some surviving dwarf galaxies (e.g. Sculptor).
Data availability
Table 6 with information on the line list used in this work is available on Zenodo at https://doi.org/10.5281/zenodo.13907515.